Shift report 20000821 08:00-16:00

From: http account (http@pii3.brahms.bnl.gov)
Date: Mon Aug 21 2000 - 16:19:20 EDT

  • Next message: http account: "Shift report 20000822 16:00-24:00"

    ------------------------------------------------------------
    Date: Mon Aug 21 16:19:20 2000, Shift: 08:00-16:00
    Supervisor: KO, with: KO & CE
    Log book#4 pages 135 - 136
    
    Shift Summary:
    ------------------------------------------------------------
    1. NO Beam during shift.
    2. Access to IR most of the day.
       BBL L4032 replaces - the new one seems to work.
       T1 was repaired - broken wire was taken out.
       BFS L1458B (T3 & RICH) was taken out.
    3. No problems to report.
    ------------------------------------------------------------
    



    This archive was generated by hypermail 2b29 : Mon Aug 21 2000 - 16:22:44 EDT